A sonar fitted in a submarine sends ultrasonic waves which reach it again after reflecting from another submarine in 50 seconds.what is the distance between the two submarines?

Grade:11

1 Answers

Rinkoo Gupta
9 years ago
In fresh water, the speed of sound at 25 degree celsius is 1497 m/s.
let the distance between the submarine is x, then
Total distance=speed*time
2x=1497 * 50
2x=74850
x=74850/2=37425
Hence the required distance is 37425 m.
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty